A researcher measures job satisfaction among married, single, and divorced employees to determine whether marital status can influence job satisfaction. Based on the following description in APA format, state the value for k, N, and n.

A one-way analysis of variance showed that job satisfaction did not vary by marital status,

F(2, 21) = 1.76, p > 0.05.

k =
N =
n =
Answer #1

Solution F (2121) = 1-76 p>0.05 since degree of freedom of numeraton is 82 so degree of freedom 1 = 2 k= d.fit=2+1 K=3] degrethank you